An 8.00 uF capacitor is connected to the terminals of an AC generator with rms voltage of 1.50 x 10^2V and a frequency of 60.0 Hz. Find the capacitive reactance and the rms current in the circuit. Draw the circuit with values in it.PLZ HELP ME OUT SIR ,,, PLZZZ EXPALIB ALL THESE STUFFF IN DETAIL

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Sohan
As we know
Xc = 1/(2 * pi* f *c) = 1/(2 * 3.14 * 60* 8 * 10^-6) = 332 ohm
Now
I rms = Delta Vrms/ Xc = 1.5 * 10² /332 = 0.452 amp
 
Regards
Arun (askIITians forum expert)